OpenStudy (anonymous):

f(x)=1/x−5 and g(x)=x+(2/x). what is f of g what is g of f

OpenStudy (lalaly):

\[f(g(x))=\frac{1}{(x+\frac{2}{x})-5}\]

OpenStudy (anonymous):

Suppose that f(x)=sqrt (x+8), g(x)=x^2+7,and h(x)=x+9. what is f of g of h

OpenStudy (lalaly):

\[g(f(x))=\frac{1}{x-5}+\frac{2}{\frac{1}{x-5}}\]

OpenStudy (lalaly):

for the second question u do the same, find g(h) first then f(g(h))